OpenStudy (anonymous):

find the domain of f(x)=[sqrt(x+2)/log10 (1-x) ] +sin^-1 x

OpenStudy (anonymous):

\[f(x)= \left[\begin{matrix}\sqrt{x+2} \\ \log_{10}(1-x) \end{matrix}\right] +\sin^{-1} x\]